OpenStudy (anonymous):

(sqrt) - 121 Simpilfy

OpenStudy (anonymous):

Havent u understood yet?

OpenStudy (anonymous):

can someone helpme out on http://openstudy.com/study#/updates/4fdb53aee4b0f2662fd17794

OpenStudy (anonymous):

No

OpenStudy (anonymous):

when your done angela will you look at this

OpenStudy (anonymous):

An urn contains pink and black balls. Four balls are randomly drawn from the urn in succession, with replacement. That is, after each draw, the selected ball is returned to the urn. What is the probability that all balls drawn from the urn are pink? Round your answer to three decimal places.

OpenStudy (anonymous):

See whts the root of 121?

OpenStudy (anonymous):

is this the question? \[\sqrt{-121}\]

OpenStudy (anonymous):

Yes

mathslover (mathslover):

\[\huge{\sqrt{-121}=\sqrt{-1*121}}\] \[\huge{\sqrt{-1}*\sqrt{121}=i(11)}\] \[\huge{11i}\]

OpenStudy (anonymous):

this is the rule you want: \[\sqrt{ab} = \sqrt{a}\sqrt{b}\] \[\sqrt{-121} = \sqrt{-1 \times 121} = \sqrt{-1}\sqrt{121} = 11i\] looks like i was beaten to it xD

OpenStudy (jiteshmeghwal9):

very good @mathslover :) deserve a medal

mathslover (mathslover):

thanks @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

:-)

OpenStudy (anonymous):

An urn contains pink and black balls. Four balls are randomly drawn from the urn in succession, with replacement. That is, after each draw, the selected ball is returned to the urn. What is the probability that all balls drawn from the urn are pink? Round your answer to three decimal places.

OpenStudy (anonymous):

9pik 7blk

OpenStudy (anonymous):

@yeababy can you try posting your own question?

OpenStudy (anonymous):

in the "ask a question" box

OpenStudy (anonymous):

no because no one wants to answer it iam forced to do this and bump like crazy

OpenStudy (anonymous):

and still no help

OpenStudy (anonymous):

so you have posted it already? link me the question

OpenStudy (anonymous):

i'll help